OpenStudy (anonymous):

The formula to find r is (A/pi)^1/2 If A=54cm^2 and pi=22/7, what is r?

OpenStudy (anonymous):

Here to make it more clear: \[\left(\begin{matrix}A \\ \pi\end{matrix}\right)^.5\] A = 54 centimeters^2 and pi = 22/7 Find r

OpenStudy (anonymous):

I keep getting 30.46010 but that is not one of my multiple choice options

geerky42 (geerky42):

Can you show work? I don't see how you got 30.460...?

OpenStudy (anonymous):

yeah (54^2/(22/7))^1/2 (2961/(22/7)) 2961/3.142 ^1/2 942.393^1/2 r = 30.698 omg what the heck? lol I got a slightly different answer that time -_-

OpenStudy (anonymous):

@geerky42 did youget my message

geerky42 (geerky42):

"54^2" Oh that's why you got wrong answer. You don't square 54, it's just that unit is squared \(cm^2\)

geerky42 (geerky42):

Should be (54/(22/7))^1/2

OpenStudy (anonymous):

@geerky42 really? :O oh let me try that

geerky42 (geerky42):

Yep. \(54~cm^2\neq54^2~cm^2\) or something

OpenStudy (anonymous):

ok I got 4.15 but that isn't one of my multiple choice options!

geerky42 (geerky42):

Really? what are options?

OpenStudy (anonymous):

my options are 7 7^(1/2) 14

OpenStudy (anonymous):

@geerky42 do you want me to show you the screenshot?

geerky42 (geerky42):

Are you sure? Because none of them are correct?

geerky42 (geerky42):

Sure

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

its a 2-part question the first part was asking for the formula to find r given this formula: A=pir^2 So maybe i got that first part wrong

geerky42 (geerky42):

It is correct that \(r = \sqrt{\dfrac{A}{\pi}}\)

OpenStudy (anonymous):

i think it is but it doesnt tell me

geerky42 (geerky42):

Should be right. Is there any report button or something? Or can you contract teacher to double check this question?

OpenStudy (anonymous):

no theres not, but this isnt the fist time I found an error in one of their tests! i'll try contacting one of the teachers

geerky42 (geerky42):

Okay, I would recommend that.